### v2

# Biomedical Open Source Software: Crucial Packages and Hidden Heroes

Abstract
Despite the importance of scientific software for research, it is often not formally recognized and rewarded. This is especially true for foundation libraries, which are used by the software packages visible to the users, being “hidden” themselves. Research stakeholders like funders, infrastructure providers and other organizations need to understand the complex network of computer programs that the modern research relies upon.
In this work, we used the CZ Software Mentions Dataset to map the dependencies of the software used in biomedical papers and find the packages critical to the software ecosystems. We propose centrality metrics for the network of software dependencies, analyze three ecosystems (PyPi, CRAN, Bioconductor) and determine the packages with the highest centrality.
1 Introduction
Since the second half of the last century, a computer has become as ubiquitous a tool of a scientific lab as balance scales and Bunsen burner in the previous ages. As a consequence, computer software is now crucial to research, bringing new methods and new scale, while offering new potential for reproducibility and extension. This is true not only for the sciences, but for scholarship more broadly (e.g., sociology and the humanities), making the software revolution both wide and deep. Yet we have very limited insight into the software actually used in research. This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and to direct science funding in a well-informed manner.
This problem is exacerbated by the fact that today, software is a complex ecosystem with a myriad of dependencies. When a researcher uses a program, they rely on the work of the people that wrote, tested and maintained it—as well as the people that wrote, tested and maintained the packages and libraries that the program depends on, either at runtime or during the compilation phase. Thus, funders and others who want to support and accelerate science need to find the critically important pieces, both among the user-facing programs and those that are not known to the end users, i.e., the former’s dependencies. This complex nature of the network of dependencies has a number of implications, including for security (Goodin, 2024) and computational reproducibility (Samuel and Mietchen, 2024).
Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace that software has become crucial. For example, software is infrequently and inconsistently formally cited (Howison and Herbsleb, 2011; Howison et al., 2015; Singh Chawla, 2016; Howison and Bullard, 2016; Knowles et al., 2021; Druskat et al., 2024). There have been recent efforts to extract informal citations from the full text of articles (Schindler et al., 2021; Du et al., 2021; Istrate et al., 2022b, a) and to evaluate the “importance” of software packages by looking at papers that cite them (Bogart et al., 2020). Unfortunately, publications sometimes do not mention all the software used in the course of the research.
Besides this, there is another kind of invisibility of scientific software. The programs visible to the user may rely on many other software packages (known in the software world as dependencies). While the end users may mention a package at the top of the dependencies stack, they are likely not even aware of the packages that are further below. These may be packages that the used program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may – and often do – depend on further packages. These latter packages thus become indirect, transitive dependencies of the user-facing program. All the work undertaken to develop, maintain, test, and distribute the underlying software is therefore not visible in the publication record itself, and is not included in derivatives such as citation graphs and knowledge graphs. The situation resembles the famous XKCD cartoon (Munroe, 2020), where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003”. The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs. There have been proposals to assign credit to these packages (Katz, 2014; Katz and Smith, 2014), but so far, the problem has not been solved. To implement the proposed transitive credit mechanisms, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc. One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) (Druskat et al., 2021). Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation graphs (Druskat, 2020). At present, the situation is quite different: sometimes even the maintainers of the down-stack software are not aware of the top stack programs that depend on their work. They need this knowledge when making breaking changes to their packages, which might negatively influence the software that depends on them (Bogart et al., 2021).
By classifying software packages as visible to the end users or primarily important for other packages, we follow the ideas described by Donald E. Stokes in his famous book (Stokes, 1997). There, Stokes distinguishes between applied research with the results visible to the general public, and the pure research, which is less visible, but important for the applied research. Following his ideas, we can put the software packages into a two dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 1). The majority of the packages will occupy the lower left corner of the plot, having a fair usage both by the authors of scientific papers and the authors of the software package. These groups may intersect. Popular packages are used by many authors. We are interested in the “Nebraska” packages, not very visible, but of critical importance. Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
Figure 1: Classification of software packages inspired by (Stokes, 1997)
We had a fortunate opportunity to explore these ideas due to the generosity of the Chan Zuckerberg Initiative (CZI). On 24–27 October 2023, CZI hosted a hackathon on Mapping the Impact of Research Software in Science (see https://github.com/chanzuckerberg/software-impact-hackathon-2023). One of the projects at this hackathon was Tracing the dependencies of open source software mentioned in the biomedical literature (Brown et al., 2023). In this project, we explored the dependencies of the open source software packages mentioned in the CZ Software Mentions dataset (Istrate et al., 2022b). We decided to limit the study to open source packages: first, because the dependencies of closed source packages are not public, and second, because we believe in the importance of open source for open and reproducible science. The result of this project is this paper. Here, we explore approaches to making the software package infrastructure underlying science more visible. We examine our findings to develop questions to better understand the idea of criticality and opportunities for improving science through adjustments to the research software ecosystems.
2 Materials and Methods
2.1 Network construction
We combined two datasets: The CZ Software Mentions dataset (Istrate et al., 2022b) (which was built using a model trained on the Softcite Dataset gold standard dataset (Du et al., 2021)), and the Ecosyste.ms software dependency dataset (Nesbitt, 2023), which is built by gathering and normalizing the dependency information from three package managers, one for Python (PyPI) and two for R (CRAN and Bioconductor). The authorship data was scraped from Github metadata, for want of consistently available authoritative software authorship metadata.
The CZ Software Mentions dataset (Istrate et al., 2022b, a) identifies which mentions were traced to which ecosystem. In this way, package names from CRAN, Bioconductor and PyPi that have been parsed by the software mentions project are collected. We used the latest release of the package at time of data collection (October 2023) along with the names of their dependencies. Those dependency names have then also been recursively fetched using the most recent release and dependencies until the full list of transitive dependencies was obtained. Limitations of this dependency resolution approach are discussed below.
After this data processing, we produced a two-mode network, with nodes for papers and nodes for packages. Edges from papers to packages were added when a package was mentioned in the full-text of the paper. Edges between packages were added when package manager descriptions indicated a dependency (including both required and suggested dependencies, and including packaging related dependencies such as those needed for development, as well as testing and building documentation). The full code for the processing is available at (Brown et al., 2023). The network is available at (Brown, 2023) in GEXF format (GEXF Working Group, 2009). It has four classes of nodes:
paper:
papers from the CZ Software Mentions dataset. They are identified by their Digital Object identifier (DOI). To estimate the impact of the papers, we separately downloaded their citation numbers from OpenAlex (Priem et al., 2022) as of November 2023 (a copy is available in the data subdirectory of (Brown et al., 2023)).
pypi, cran, bioconductor:
Software from the corresponding ecosystems. We used CZI ID ((Istrate et al., 2022b, a) as the identifier. We did not attempt to identify the same software across the ecosystems (see the discussion below).
The edges are directed and weighted. An edge from software node $A$ to software node $B$ means that software $A$ depends on software $B$ as determined by the corresponding package manager. An edge from a paper node to a software node means that the given software is mentioned in the paper, the weight corresponds to the number of citations the paper received. In what follows, we will consider both the weighted and unweighted version of this network, since the weights have pros and cons, as discussed below.
2.2 Network analysis
To analyze the network, we relied on directed centrality analysis (Borgatti and Everett, 2021). There are several possible options for centrality measures. Our choice was determined by the following considerations. First, we wanted a centrality metric that can account for papers even though they have no incoming edges in our network (in other words, papers do not receive centrality but should contribute to the centrality of software packages). This criterion excludes eigenvector centrality (Bonacich, 2007). Second, we wanted a centrality metric that can account for the weight of edges from papers to software packages (i.e., some papers contribute more centrality through their citation counts). This criterion excludes the PageRank algorithm (Brin and Page, 1998) which normalizes weights of out-degree. We found that Katz centrality (Katz, 1953) satisfies all our criteria. This approach gives us the attenuation factor $\beta$ as a free parameter to control the importance of papers (which we set equal to 1), such that paper nodes contribute a factor proportional to $\beta$ times their citation count to the software packages they mention. In turn, these software packages contribute a factor proportional to their own centrality to their dependencies. A package can therefore be central by receiving mentions from well-cited papers, by having central dependents, or by a combination of both.
Refer to caption (a)
Refer to caption (b)
Figure 2: (a) Network visualization of software packages from three ecosystems (from CRAN in green, PyPI in blue, and Bioconductor in pink) connected through their dependencies within their ecosystem and interconnected through papers that mention them. We label the top 3 most central packages in each ecosystem: ggplot2, SAM, and PRISMA for CRAN, velvet, tophat and pymol for PyPI and DeSeq2, edgeR and limma for Bioconductor. The core of the network is dominated by CRAN and PyPI dependencies, despite the fact that three of the five most central packages come from Bioconductor. (b) The top part of the above network, with papers added (in grey) to illustrate how PRISMA can be central due to many mentions in papers.
3 Results
In Figure 2, we show the overall network of software packages connected through their dependencies within their ecosystem and interconnected through the papers that mention them. Edges from papers to software packages are directed and weighted by the number of citations the paper received. Edges between software dependencies are directed from a dependent to a dependency.
We found a dense core of popular packages that receive many mentions (e.g. ggplot2 in CRAN, tophat in PyPI and limma in Bioconductor), some of which have many dependencies themselves (e.g. ggplot2). We also found less central packages mentioned heavily by specific communities (e.g. PRISMA in CRAN or pymol in PyPI).
Analyzing the entire ecosystems of software dependency, we found that roughly 10% of software packages are part of dependency loops (i.e., cycles in the dependency networks). Interestingly, we found no cycles whatsoever in the connected component of networks that have received software mentions, and their dependencies. We discuss this finding below.
In Figure 3, we show the distribution of packages from the three investigated ecosystems over Katz centrality and mention counts. We used three different ways to calculate centrality: (a) unweighted graph, (b) weighted complete graph, and (c) the largest connected component of the weighted graph. It is remarkable that all three models give rather similar results for many packages. In the next section, we discuss the insights from the cases when the different methods do not agree. As expected, we found a dense cluster of packages with low Katz centrality and low mention counts, i.e., the “Majority” quadrant (see Figure 1).
Figure 3: Distribution of packages by Katz centrality and counts of their mentions in papers. Katz centrality is calculated for an unweighted graph, for a weighted graph with all nodes, or just for the largest connected cluster (LCC). In the calculations, we assumed $\beta=1$.
4 Discussion
The absence of cycles in the network of open-source software packages used in our sample of biomedical science is quite interesting. It suggests a more robust design structure in the universe of scientific software than in the general software world. The lack of cycles also improves our analysis with Katz centrality, as it excludes feedback mechanisms that can artificially inflate the centrality of packages on a loop. Packages with low mentions and high Katz centrality are therefore even more critical, as they represent essential dependencies of packages that enable large volumes of research without receiving direct attention. These are exactly the “Nebraskan” packages referenced by Munroe (2020) that we set out to find. Figure 3 seems to show one clear example of such a package: velvet, a PyPI package with a mention count of $197$, but a Katz centrality value calculated as $0.105$ for the largest connected cluster and as $0.053$ for the complete graph. Interestingly, for the unweighted representation, this package has a low centrality of $0.007$.
Generally, the results verify our intuition that the majority of software packages will have a low Katz centrality and relatively few mentions. The outliers are then the most interesting packages, because they are central and often mentioned (“Pasteur”), or because their centrality outweighs their mentions (“Nebraska”), or vice versa (popular packages).
The difference between the unweighted and weighted version of our analysis, in Figure 3 can help us parse different subtypes of these outliers. For example, the velvet package discussed above is central solely because it is mentioned by one highly cited paper (and dozens of less cited ones). This strategy is not central in the unweighted version of the analysis, since citation counts are ignored. Central packages under this representation therefore also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
Altogether, our paper tries to quantify impact as importance, and to put some numbers behind the intuition. Of course, this is by itself a perilous process: there could be various flavors of both impact and importance, and people may reasonably disagree about the details. Our contribution is to propose a certain quantification, which agrees with our intuitive understanding.
Besides the possible problems with the exact definition of importance, our work has a number of limitations:
  1. 1.
    The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions Dataset (see the discussion in (Istrate et al., 2022a)). Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms (Druskat et al., 2024). Notably, this affects the above-mentioned “Nebraskan” package velvet, which is linked to a Python package of the same name that provides signal processing and communications algorithms in Python (available on PyPI [^1] https://pypi.org/project/velvet/). Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short read assembly using de Bruijn graphs in genomics, as described in (Zerbino and Birney, 2008) and implemented mostly in C (Zerbino et al., 2014). Sometimes an open source package shares a name with a proprietary software (e. g. PRISM, PACE), which leads to confusion and incorrect usage counts.
  2. 2.
    Our approach is limited to the open source packages included into one of the chosen package managers. This excludes proprietary software like Excel or GraphPad Prism, and open source non-packaged software like Gephi (Bastian et al., 2009). The number of missing packages may be estimated from Table 14 of (Istrate et al., 2022a) which gives the coverage by GitHub at 64.39%, the coverage by CRAN at 8.36%, the coverage by PyPI at 5.86%, and the coverage by Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages – perhaps growing in number – are not hosted by GitHub, their fraction remains anecdotally low. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e. between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open source non-package software, it may be feasible to identify source repository URLs and then to identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services); nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers222as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies.
  3. 3.
    In our version of the dependency chart, we always used the latest version of any package. In reality, dependencies change between the versions. A study of development logs often reveals messages like “deleted the dependence on XYZ” or “added the dependence on ABC”. These changes often make the work of package management software very difficult. A software dependency graph is a living network, constantly evolving, with links added and deleted. In this study, we effectively collapse the time, making an (imperfect) snapshot of the long movie.
    A future improvement would be to use each package ecosystem’s specific dependency resolution algorithm to compute the full transitive dependency tree for each mentioned software package. An even further extension would be to attempt to do this for the version of the package most likely used by a specific publication, based on aligning publication date with the current version used at that time.
  4. 4.
    Our method captures the off the shelf software packages used for the papers. It does not capture ad hoc software written specifically for the given paper, which may be contained in the code accompanying the papers. This code may load software libraries, and thus add software dependencies absent in our graph.
  5. 5.
    The dependencies reflected in the package managers do not necessarily reflect the actual usage of the software. A library may be loaded, but not used for the actual execution path used in the given paper. A more reliable method would be to reproduce the analysis done in each paper and capture the actual library calls—which is probably prohibitively difficult for the number of papers covered.
  6. 6.
    Many Python and R packages ultimately depend on libraries outside their ecosystems (most often C and Fortran libraries). These dependencies are not easily captured by the package management software, and thus are outside of our analysis.
  7. 7.
    Our dependency graph does not include the concept of alternatives, when package $A$ depends on the functionality that can be provided either by package $B_{1}$, or package $B_{2}$. Some package managers like Debian’s apt have the possibility of specifying alternatives, but this is not a common feature of package managers.
  8. 8.
    Different communities place different emphasis on infrastructure elements such as tests. This may lead to the over-representation of testing infrastructure packages in some cases, for example in the Python ecosystem. We do not try to attempt to decide which elements of the infrastructure are “actually important”, instead relying on the judgment of the community of users and developers—while recognizing that such judgment may be different across the landscape.
  9. 9.
    There are several different Python package managers with greatly overlapping dependency graphs. We used PyPI for this study. Some alternatives like pip and Conda are also worth investigating.
5 Conclusions and future work
In this work, we investigated the network of software packages used in the biomedical papers available through PubMed Central. We demonstrated that these packages follow the structure of “Stokes’ diagram” (famous for ”Pasteur’s quadrant”), with some packages highly visible to the end users, and some packages less visible, but important in the network due to their dependencies. We discussed the use of Katz centrality in discovering the important packages and found examples of such packages in the biomedical field. We also discovered that the structure of the network of the software packages used in biomedical science is different from the structure of generic software packages: the absence of dependence loops may indicate a more robust, or intentional, design.
These findings and insights as well as the underlying methodology can be used in funding and development of open source scientific software.
Of course, there are many ways to extend this work. It would be interesting, for instance, to analyze common workflows for different disciplines, perhaps using co-occurrences of mentions, and map them into the dependency graph. This might help to discover packages important for specific sub-fields of biomedical sciences. Similarly, adding temporal dependencies to our graph may help to discover and predict development trends. Also, individual dependencies may differ in the actual impact that they have on the dependent software. Developing and applying a metric that describes these differences in impact would make it possible to refine our approach to quantifying the importance of research software dependencies.
Acknowledgments
The authors are grateful to the Chan Zuckerberg Initiative for making this research possible and for sponsoring the Mapping the Impact of Research Software in Science hackathon. L.H.-D. acknowledges support from the National Institute of General Medical Sciences under the 2P20GM125498 Centers of Biomedical Research Excellence Award and from Google Open Source through the Open-Source Complex Ecosystems And Networks (OCEAN) project. D.M. acknowledges support from the German Research Council (DFG) through the MaRDI project (460135501) as well as through the BASE4NFDI/KGI4NFDI project (521453681). J.F.P. acknowledges support from CNPq and FAPERJ (E-26/210.478/2024).
Data and code availability
The code for this paper is available at (Brown et al., 2023), and the data are available at (Brown, 2023).

### v3

@@
 Despite the importance of scientific software for research, it is often not formally recognized and rewarded.
-This is especially true for foundation libraries, which are used by the software packages visible to the users, being “hidden” themselves.
-Research stakeholders like funders, infrastructure providers and other organizations need to understand the complex network of computer programs that the modern research relies upon.
-In this work, we used the CZ Software Mentions Dataset to map the dependencies of the software used in biomedical papers and find the packages critical to the software ecosystems.
-We propose centrality metrics for the network of software dependencies, analyze three ecosystems (PyPi, CRAN, Bioconductor) and determine the packages with the highest centrality.
+This is especially true for foundational libraries, which are hidden below packages visible to the users (and thus doubly hidden, since even the packages directly used in research are frequently not visible in the paper).
+Research stakeholders like funders, infrastructure providers, and other organizations need to understand the complex network of computer programs that contemporary research relies upon.
+In this work, we use the CZ Software Mentions Dataset to map the upstream dependencies of software used in biomedical papers and find the packages critical to scientific software ecosystems.
+We propose centrality metrics for the network of software dependencies, analyze three ecosystems (PyPi, CRAN, Bioconductor), and determine the packages with the highest centrality.
+Author summary
+Scientists today use software as an important tool for their research.
+The progress of science depends on the people who write software being properly recognized, rewarded, and incentivised.
+This means that those who fund research or promote scientists need to know the impact of software and identify the most important programs.
+The universe of scientific software is complex: some people write programs used by researchers, while some people write programs used by other programs.
+The latter are often even less visible than the former, but they are important for the progress of science.
+In this work, we establish metrics to measure the impact of both types of software.
+We introduce the network of dependencies (program A used by researchers depends on programs B, C, and D) and use it to calculate the impact of a large number of packages used in the biomedical literature.
+In doing so, we identify important software packages, regardless of their direct visibility to end users.
+We conclude with a discussion of the limitations of our methodology and approaches to improve network validity.
 1 Introduction
-Since the second half of the last century, a computer has become as ubiquitous a tool of a scientific lab as balance scales and Bunsen burner in the previous ages.
+Since the second half of the last century, a computer has become as ubiquitous a tool of a scientific lab as balance scales and a Bunsen burner were in previous ages.
 As a consequence, computer software is now crucial to research, bringing new methods and new scale, while offering new potential for reproducibility and extension.
-This is true not only for the sciences, but for scholarship more broadly (e.g., sociology and the humanities), making the software revolution both wide and deep.
+This is true not only for the natural sciences and mathematics but for scholarship more broadly (e.g., sociology and the humanities), making the software revolution both wide and deep.
 Yet we have very limited insight into the software actually used in research.
-This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and to direct science funding in a well-informed manner.
-This problem is exacerbated by the fact that today, software is a complex ecosystem with a myriad of dependencies.
-When a researcher uses a program, they rely on the work of the people that wrote, tested and maintained it—as well as the people that wrote, tested and maintained the packages and libraries that the program depends on, either at runtime or during the compilation phase.
-Thus, funders and others who want to support and accelerate science need to find the critically important pieces, both among the user-facing programs and those that are not known to the end users, i.e., the former’s dependencies.
-This complex nature of the network of dependencies has a number of implications, including for security (Goodin, 2024) and computational reproducibility (Samuel and Mietchen, 2024).
-Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace that software has become crucial.
-For example, software is infrequently and inconsistently formally cited (Howison and Herbsleb, 2011; Howison et al., 2015; Singh Chawla, 2016; Howison and Bullard, 2016; Knowles et al., 2021; Druskat et al., 2024).
-There have been recent efforts to extract informal citations from the full text of articles (Schindler et al., 2021; Du et al., 2021; Istrate et al., 2022b, a) and to evaluate the “importance” of software packages by looking at papers that cite them (Bogart et al., 2020).
-Unfortunately, publications sometimes do not mention all the software used in the course of the research.
+This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and direct science funding in a well-informed manner.
+Fig 1: XKCD on software dependency and ecosystem fragility [1].
+Also has roll-over text of Someday ImageMagick will finally break for good and we’ll have a long period of scrambling as we try to reassemble civilization from the rubble.
+Used with permission: https://creativecommons.org/licenses/by-nc/2.5/
+Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace as software has become crucial.
+For example, software is infrequently and inconsistently formally cited [2, 3, 4, 5, 6, 7].
+There have been recent efforts to extract informal citations from the full text of articles [8, 9, 10, 11] and to evaluate the “importance” of software packages by looking at papers that cite them [12].
+Unfortunately, publications sometimes do not mention all the software used in the course of research.
 Besides this, there is another kind of invisibility of scientific software.
-The programs visible to the user may rely on many other software packages (known in the software world as dependencies).
-While the end users may mention a package at the top of the dependencies stack, they are likely not even aware of the packages that are further below.
-These may be packages that the used program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may – and often do – depend on further packages.
+The programs visible to the end user may rely on many other software packages (known in the software world as dependencies).
+While the end users may mention a package at the top of the dependency stack, they are likely not even aware of the full set of packages that are further below.
+These may be packages that the utilized program depends on directly (direct dependencies), or indirectly, where the direct dependencies in turn may depend on further packages.
 These latter packages thus become indirect, transitive dependencies of the user-facing program.
-All the work undertaken to develop, maintain, test, and distribute the underlying software is therefore not visible in the publication record itself, and is not included in derivatives such as citation graphs and knowledge graphs.
-The situation resembles the famous XKCD cartoon (Munroe, 2020), where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003”.
+This complex nature of the network of dependencies has a number of implications, including for security [13] and computational reproducibility [14].
+In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
+The situation resembles the famous XKCD cartoon shown in  Figure 1, where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1].
 The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs.
-There have been proposals to assign credit to these packages (Katz, 2014; Katz and Smith, 2014), but so far, the problem has not been solved.
-To implement the proposed transitive credit mechanisms, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc.
-One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) (Druskat et al., 2021).
-Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation graphs (Druskat, 2020).
-At present, the situation is quite different: sometimes even the maintainers of the down-stack software are not aware of the top stack programs that depend on their work.
-They need this knowledge when making breaking changes to their packages, which might negatively influence the software that depends on them (Bogart et al., 2021).
+The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
+There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termedtransitive credit [15, 16] but so far the problem has not yet been solved.
+To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc.
+One way for projects to do this is the inclusion of citation information for their own software outputs as well as for their direct dependencies, e.g., a citation file in the Citation File Format (CFF) [17].
+Were such citation information available for the complete dependency stack of a program, transitive credit could be implemented by building weighted software citation networks [18].
+At present, though, the situation is quite different: sometimes even the maintainers of the lower-stack software are not aware of the upper-stack programs that depend on their work.
+They need this knowledge when making breaking changes to their packages, which might negatively influence the software that depends on them [19].
 By classifying software packages as visible to the end users or primarily important for other packages, we follow the ideas described by Donald E.
-Stokes in his famous book (Stokes, 1997).
-There, Stokes distinguishes between applied research with the results visible to the general public, and the pure research, which is less visible, but important for the applied research.
-Following his ideas, we can put the software packages into a two dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 1).
-The majority of the packages will occupy the lower left corner of the plot, having a fair usage both by the authors of scientific papers and the authors of the software package.
-These groups may intersect.
+Stokes in his famous book, Pasteur’s Quadrant [20].
+There, Stokes distinguishes between applied research with the results visible to the general public, and pure research, which is less visible, but important for the applied research.
+He discussed a two-dimensional diagram of scientific works, with the importance of the works for the applications on the $x$ axis and the quest for fundamental understanding on the $y$ axis.
+Thus, the upper left quadrant is occupied by the purely fundamental research personified by Niels Bohr.
+The lower right quadrant is occupied by purely applied research personified by Thomas Edison.
+The upper right quadrant is occupied by the works that have both fundamental and practical value.
+Stokes chose Louis Pasteur as an example of such works.
+While this picture is rather simplified (Bohr’s works have significant practical value, and Edison’s experiments stimulated fundamental research), it gives important insight into scientific endeavor.
+Following these ideas, we can put the software packages into a two-dimensional plane with the axes corresponding to the frequency of software mentions and a measure of network centrality (Figure 2).
+The majority of the packages will likely occupy the lower left corner of the plot, having a small number of mentions from the authors of scientific papers and limited centrality from usage in other software packages.
 Popular packages are used by many authors.
@@
 Of course, one can think about highly visible and critical “Pasteur” packages, which are used both directly and as a foundation for other libraries.
-Figure 1: Classification of software packages inspired by (Stokes, 1997)
+Fig 2: Classification of software packages inspired by [20]
 We had a fortunate opportunity to explore these ideas due to the generosity of the Chan Zuckerberg Initiative (CZI).
 On 24–27 October 2023, CZI hosted a hackathon on Mapping the Impact of Research Software in Science (see https://github.com/chanzuckerberg/software-impact-hackathon-2023).
-One of the projects at this hackathon was Tracing the dependencies of open source software mentioned in the biomedical literature (Brown et al., 2023).
-In this project, we explored the dependencies of the open source software packages mentioned in the CZ Software Mentions dataset (Istrate et al., 2022b).
-We decided to limit the study to open source packages: first, because the dependencies of closed source packages are not public, and second, because we believe in the importance of open source for open and reproducible science.
-The result of this project is this paper.
+One of the projects at this hackathon was Tracing the dependencies of open source software mentioned in the biomedical literature [21].
+In this project, we explored the dependencies of the open source software packages mentioned in the CZ Software Mentions dataset [10].
+The dataset was created using a machine learning system trained on the SoftCite gold standard dataset of manually annotated software mentions [9] and consists of extractions from 2.4 million biomedical papers.
+The methodology and its limitations are discussed in [11].
+In particular, the work on the disambiguation of packages can be improved.
+Later, we discuss how these limitations influence our results.
+We decided to limit our study to open-source packages: first, because the dependencies of closed-source packages are not public, and second, because we believe in the importance of open source for open and reproducible science.
 Here, we explore approaches to making the software package infrastructure underlying science more visible.
@@
 2.1 Network construction
-We combined two datasets: The CZ Software Mentions dataset (Istrate et al., 2022b) (which was built using a model trained on the Softcite Dataset gold standard dataset (Du et al., 2021)), and the Ecosyste.ms software dependency dataset (Nesbitt, 2023), which is built by gathering and normalizing the dependency information from three package managers, one for Python (PyPI) and two for R (CRAN and Bioconductor).
-The authorship data was scraped from Github metadata, for want of consistently available authoritative software authorship metadata.
-The CZ Software Mentions dataset (Istrate et al., 2022b, a) identifies which mentions were traced to which ecosystem.
-In this way, package names from CRAN, Bioconductor and PyPi that have been parsed by the software mentions project are collected.
-We used the latest release of the package at time of data collection (October 2023) along with the names of their dependencies.
-Those dependency names have then also been recursively fetched using the most recent release and dependencies until the full list of transitive dependencies was obtained.
-Limitations of this dependency resolution approach are discussed below.
-After this data processing, we produced a two-mode network, with nodes for papers and nodes for packages.
-Edges from papers to packages were added when a package was mentioned in the full-text of the paper.
-Edges between packages were added when package manager descriptions indicated a dependency (including both required and suggested dependencies, and including packaging related dependencies such as those needed for development, as well as testing and building documentation).
-The full code for the processing is available at (Brown et al., 2023).
-The network is available at (Brown, 2023) in GEXF format (GEXF Working Group, 2009).
-It has four classes of nodes:
+We combined three datasets: The CZ Software Mentions dataset [10], which was constructed using a machine learning model trained on the SoftCite Gold Standard annotations [9], the Ecosyste.ms software dependency dataset [22], which is built by gathering and normalizing the dependency information from multiple software ecosystems, including one for the Python programming language (PyPI) and two for the R programming language (CRAN and Bioconductor), and finally we used OpenAlex [23] for information on how frequently papers had been cited.
+The software authorship data were scraped from GitHub metadata.
+The CZ Software Mentions dataset [10, 11] includes mentions from 2.4 million biomedical papers and identifies which mentions were traced to which ecosystem.
+In this way, package names from CRAN, Bioconductor, and PyPI are parsed by the software mentions extraction model, collected, and finally disambiguated.
+We collected information for each package, including its dependencies, using the latest release of the package at the time of data collection (October 2023).
+Dependencies were then recursively retrieved using the most recent release and following dependencies until the full list of transitive dependencies was obtained.
+Limitations of our dependency resolution approach are discussed later.
+After this data processing, we produced a two-mode network, with nodes for papers and nodes for software packages.
+Edges from papers to software packages were added when a package was mentioned in the full text of the article.
+Edges between packages were added when metadata descriptions indicated a required dependency.
+The full code for the processing is available at [21].
+The network is available at [24] in GEXF format [25] and has four classes of nodes:
 paper:
@@
 They are identified by their Digital Object Identifier (DOI).
-To estimate the impact of the papers, we separately downloaded their citation numbers from OpenAlex (Priem et al., 2022) as of November 2023 (a copy is available in the data subdirectory of (Brown et al., 2023)).
+To estimate the impact of the papers, we separately downloaded their citation numbers from OpenAlex [23] as of November 2023 (a copy is available in the data subdirectory of [21]).
 pypi, cran, bioconductor:
-Software from the corresponding ecosystems.
-We used CZI ID ((Istrate et al., 2022b, a) as the identifier.
-We did not attempt to identify the same software across the ecosystems (see the discussion below).
+Software from the corresponding ecosystems. We used CZI ID ([10, 11] as the identifier. We did not attempt to identify the same software across the ecosystems (see the discussion below).
 The edges are directed and weighted.
-An edge from software node $A$ to software node $B$ means that software $A$ depends on software $B$ as determined by the corresponding package manager.
-An edge from a paper node to a software node means that the given software is mentioned in the paper, the weight corresponds to the number of citations the paper received.
-In what follows, we will consider both the weighted and unweighted version of this network, since the weights have pros and cons, as discussed below.
+An edge from one software node $A$ to another software node $B$ means that software $A$ depends on software $B$ as determined by the corresponding metadata.
+An edge from a paper node to a software node means that the given software is mentioned in the paper, and the weight corresponds to the number of citations the paper received.
+In what follows, we will consider both the weighted and unweighted versions of this network, as utilization of the weights has both positive and negative aspects to consider during analysis and interpretation.
 2.2 Network analysis
-To analyze the network, we relied on directed centrality analysis (Borgatti and Everett, 2021).
+To analyze the network, we relied on directed centrality analysis [26].
 There are several possible options for centrality measures.
 Our choice was determined by the following considerations.
-First, we wanted a centrality metric that can account for papers even though they have no incoming edges in our network (in other words, papers do not receive centrality but should contribute to the centrality of software packages).
-This criterion excludes eigenvector centrality (Bonacich, 2007).
-Second, we wanted a centrality metric that can account for the weight of edges from papers to software packages (i.e., some papers contribute more centrality through their citation counts).
-This criterion excludes the PageRank algorithm (Brin and Page, 1998) which normalizes weights of out-degree.
-We found that Katz centrality (Katz, 1953) satisfies all our criteria.
-This approach gives us the attenuation factor $\beta$ as a free parameter to control the importance of papers (which we set equal to 1), such that paper nodes contribute a factor proportional to $\beta$ times their citation count to the software packages they mention.
-In turn, these software packages contribute a factor proportional to their own centrality to their dependencies.
+First, we wanted a centrality metric that can account for papers even if they have no incoming edges in our network (in other words, papers do not receive centrality but should contribute to the centrality of software packages).
+This criterion excludes eigenvector centrality [27].
+Second, we wanted a centrality metric that can account for the weight of edges from papers to software packages (i.e., some papers contribute more centrality because they are more highly cited.).
+This criterion excludes the PageRank algorithm [28], which normalizes weights of out-degree.
+We found that Katz centrality [29] satisfies all our criteria.
+Katz centrality can be interpreted as the importance of the node for the diffusion processes on the network.
+This corresponds to the intuition that innovation is a diffusion-like process [30].
+Katz centrality gives us the attenuation factor $\beta$ as a free parameter to control the importance of papers (which we set equal to 1), such that paper nodes contribute a factor proportional to $\beta$ times their citation count to the software packages they mention (we follow the formulation of Katz centrality in [31]).
+In turn, software packages contribute a factor proportional of their own centrality to their dependencies.
 A package can therefore be central by receiving mentions from well-cited papers, by having central dependents, or by a combination of both.
@@
 Refer to caption (b)
-Figure 2: (a) Network visualization of software packages from three ecosystems (from CRAN in green, PyPI in blue, and Bioconductor in pink) connected through their dependencies within their ecosystem and interconnected through papers that mention them.
+Fig 3: (a) Network visualization of software packages from three ecosystems (from CRAN in green, PyPI in blue, and Bioconductor in pink) connected through their dependencies within their ecosystem and interconnected through papers that mention them.
 We label the top 3 most central packages in each ecosystem: ggplot2, SAM, and PRISMA for CRAN, velvet, tophat and pymol for PyPI and DeSeq2, edgeR and limma for Bioconductor.
@@
 3 Results
-In Figure 2, we show the overall network of software packages connected through their dependencies within their ecosystem and interconnected through the papers that mention them.
+In Figure 3, we show the overall network of software packages connected by their dependencies within each ecosystem and interconnected through the papers that mention them.
 Edges from papers to software packages are directed and weighted by the number of citations the paper received.
@@
 ggplot2).
-We also found less central packages mentioned heavily by specific communities (e.g.
+We also found packages specific to certain communities (e.g.
 PRISMA in CRAN or pymol in PyPI).
-Analyzing the entire ecosystems of software dependency, we found that roughly 10% of software packages are part of dependency loops (i.e., cycles in the dependency networks).
-Interestingly, we found no cycles whatsoever in the connected component of networks that have received software mentions, and their dependencies.
+Analyzing the entire ecosystem of software dependencies, we found that roughly 10% of software packages are part of dependency loops (i.e., cycles in the dependency networks).
+Interestingly, we found no cycles in the connected components of networks that have received software mentions, and their dependencies.
 We discuss this finding below.
+Note that contemporary package managers usually resolve the dependency loops by installing all packages in the loop simultaneously, so the loops, while indicating rather sloppy software practices, are not fatal.
 In Figure 4, we show the distribution of packages from the three investigated ecosystems over Katz centrality and mention counts.
-We used three different ways to calculate centrality: (a) unweighted graph, (b) weighted complete graph, and (c) the largest connected component of the weighted graph.
+We used three different ways to calculate centrality: (a) unweighted graph, (b) weighted complete graph, and (c) the largest connected components of the weighted graph (calculated separately for each ecosystem).
 It is remarkable that all three models give rather similar results for many packages.
@@
 As expected, we found a dense cluster of packages with low Katz centrality and low mention counts, i.e., the “Majority” quadrant (see Figure 2).
-Figure 3: Distribution of packages by Katz centrality and counts of their mentions in papers.
-Katz centrality is calculated for an unweighted graph, for a weighted graph with all nodes, or just for the largest connected cluster (LCC).
+Table 1: Summary statistics for the number of mentions in papers.
+“Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Gini” is Gini coefficient for the mentions count.
+Ecosystem Number Dependency only Mentions count
+Median IQR Max Gini
+bioconductor $1018$ 0.047 151 277 013 752 50.04715127701375250.047\,151\,277\,013\,752\,50.047 151 277 013 752 5 $24$ $68$ 21 2572125721\,25721 257 0.843 712 143 943 3090.8437121439433090.843\,712\,143\,943\,3090.843 712 143 943 309
+cran $3594$ 0.113 244 296 048 970.113244296048970.113\,244\,296\,048\,970.113 244 296 048 97 $9$ $34$ 21 9602196021\,96021 960 0.889 196 914 901 7220.8891969149017220.889\,196\,914\,901\,7220.889 196 914 901 722
+pypi $5596$ 0.143 495 353 824 160.143495353824160.143\,495\,353\,824\,160.143 495 353 824 16 $4$ $17$ 20 3222032220\,32220 322 0.888 156 736 619 3960.8881567366193960.888\,156\,736\,619\,3960.888 156 736 619 396
+Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, “Gini” is Gini coefficient.
+Ecosystem Median IQR Max Gini
+Unweighted graph
+bioconductor 0.001 263 667 285 359 560.001263667285359560.001\,263\,667\,285\,359\,560.001 263 667 285 359 56 0.000 651 274 677 839 160.000651274677839160.000\,651\,274\,677\,839\,160.000 651 274 677 839 16 0.217 934 004 136 6260.2179340041366260.217\,934\,004\,136\,6260.217 934 004 136 626 0.549 779 090 129 2210.5497790901292210.549\,779\,090\,129\,2210.549 779 090 129 221
+cran 0.001 166 462 109 562 670.001166462109562670.001\,166\,462\,109\,562\,670.001 166 462 109 562 67 0.000 486 025 878 984 450.000486025878984450.000\,486\,025\,878\,984\,450.000 486 025 878 984 45 0.277 617 982 075 9160.2776179820759160.277\,617\,982\,075\,9160.277 617 982 075 916 0.496 401 393 691 9490.4964013936919490.496\,401\,393\,691\,9490.496 401 393 691 949
+pypi 0.001 080 921 554 861 410.001080921554861410.001\,080\,921\,554\,861\,410.001 080 921 554 861 41 0.000 194 410 351 593 780.000194410351593780.000\,194\,410\,351\,593\,780.000 194 410 351 593 78 0.191 008 170 440 8880.1910081704408880.191\,008\,170\,440\,8880.191 008 170 440 888 0.351 314 534 735 6710.3513145347356710.351\,314\,534\,735\,6710.351 314 534 735 671
+Weighted, complete graph
+bioconductor 0.001 431 553 182 867 530.001431553182867530.001\,431\,553\,182\,867\,530.001 431 553 182 867 53 0.000 322 302 811 767 7570.0003223028117677570.000\,322\,302\,811\,767\,7570.000 322 302 811 767 757 0.105 238 137 865 5360.1052381378655360.105\,238\,137\,865\,5360.105 238 137 865 536 0.358 203 102 844 1960.3582031028441960.358\,203\,102\,844\,1960.358 203 102 844 196
+cran 0.001 388 172 783 386 690.001388172783386690.001\,388\,172\,783\,386\,690.001 388 172 783 386 69 0.000 147 428 626 544 9850.0001474286265449850.000\,147\,428\,626\,544\,9850.000 147 428 626 544 985 0.128 618 817 548 2220.1286188175482220.128\,618\,817\,548\,2220.128 618 817 548 222 0.253 046 882 585 3520.2530468825853520.253\,046\,882\,585\,3520.253 046 882 585 352
+pypi 0.001 382 750 233 451 590.001382750233451590.001\,382\,750\,233\,451\,590.001 382 750 233 451 59 0.000 104 384 086 250 760.000104384086250760.000\,104\,384\,086\,250\,760.000 104 384 086 250 76 0.073 772 436 229 610.073772436229610.073\,772\,436\,229\,610.073 772 436 229 61 0.180 193 304 141 0160.1801933041410160.180\,193\,304\,141\,0160.180 193 304 141 016
+Weighted, LCC only
+bioconductor 0.002 883 015 709 935 830.002883015709935830.002\,883\,015\,709\,935\,830.002 883 015 709 935 83 0.000 759 200 923 709 920.000759200923709920.000\,759\,200\,923\,709\,920.000 759 200 923 709 92 0.210 742 476 047 3810.2107424760473810.210\,742\,476\,047\,3810.210 742 476 047 381 0.371 663 779 294 7510.3716637792947510.371\,663\,779\,294\,7510.371 663 779 294 751
+cran 0.002 793 430 476 011 270.002793430476011270.002\,793\,430\,476\,011\,270.002 793 430 476 011 27 0.000 369 917 791 154 9370.0003699177911549370.000\,369\,917\,791\,154\,9370.000 369 917 791 154 937 0.257 562 976 941 2260.2575629769412260.257\,562\,976\,941\,2260.257 562 976 941 226 0.280 390 754 327 0070.2803907543270070.280\,390\,754\,327\,0070.280 390 754 327 007
+pypi 0.002 785 286 363 836 310.002785286363836310.002\,785\,286\,363\,836\,310.002 785 286 363 836 31 0.000 312 869 642 721 380.000312869642721380.000\,312\,869\,642\,721\,380.000 312 869 642 721 38 0.147 731 480 149 7160.1477314801497160.147\,731\,480\,149\,7160.147 731 480 149 716 0.220 244 615 942 4560.2202446159424560.220\,244\,615\,942\,4560.220 244 615 942 456
+Figure 4 suggests large differences between the packages in the majority quadrant and the packages in the rest of the system.
+This is further illustrated by Tables1 and 2.
+The sum of the median and half of the interquartile range is several orders of magnitude smaller than the maximal value for both the number of mentions and centrality.
+A convenient measure of this inequality is the Gini coefficient [32].
+Gini coefficients close to one characterize a very unequal distribution, while Gini coefficients close to zero correspond to a uniform distribution.
+As seen from Table 1, the distribution of mention counts is wildly unequal with high Gini coefficients.
+The distribution of centrality (Table 2) is smoother, especially for the weighted graph, but is still rather unequal.
+Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLoS CompBio for suggesting this metric).
+It is shown in Table 1.
+In Tables 3, 4, and 5 we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
+  1. 1.
+    “Pasteur”: the highest sum of mentions and centrality percentile.
+  2. 2.
+    Popular: the highest mentions percentile while centrality percentile is less than $0.5$.
+  3. 3.
+    “Nebraska”: the highest centrality percentile while mentions percentile is less than $0.5$.
+Fig 4: Distribution of packages by Katz centrality and counts of their mentions in papers.
+Katz centrality is calculated for an unweighted graph, for a weighted graph with all nodes, or just for the largest connected cluster (LCC) for each ecosystem.
 In the calculations, we assumed $\beta=1$.
+Table 3: Software quadrants, unweighted graph
+Software Ecosystem Mentions Centrality
+Count Percentile Value Percentile
+“Pasteur”
+PRISMA cran 15 7971579715\,79715 797 $0.9997$ 0.277 620.277620.277\,620.277 62 $1$
+DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.198 880.198880.198\,880.198 88 $0.9997$
+ggplot2 cran 14 4411444114\,44114 441 $0.9995$ 0.234 070.234070.234\,070.234 07 $0.9999$
+pymol pypi 14 6041460414\,60414 604 $0.9996$ 0.191 010.191010.191\,010.191 01 $0.9996$
+PERMANOVA cran 21 9602196021\,96021 960 $1$ $0.1219$ $0.9992$
+limma bioconductor 13 4511345113\,45113 451 $0.9993$ 0.217 930.217930.217\,930.217 93 $0.9998$
+SAM cran 13 0481304813\,04813 048 $0.9992$ 0.144 250.144250.144\,250.144 25 $0.9994$
+edgeR bioconductor 12 9231292312\,92312 923 $0.9991$ 0.171 370.171370.171\,370.171 37 $0.9995$
+lme4 cran 10 5101051010\,51010 510 $0.999$ 0.131 910.131910.131\,910.131 91 $0.9993$
+WGCNA cran 13 9151391513\,91513 915 $0.9994$ 0.097 210.097210.097\,210.097 21 $0.9988$
+vegan cran 10 0121001210\,01210 012 $0.9989$ 0.121 210.121210.121\,210.121 21 $0.9991$
+tophat pypi $7642$ $0.9987$ $0.1184$ $0.999$
+STAR cran $9091$ $0.9988$ 0.110 520.110520.110\,520.110 52 $0.9989$
+Popular
+GSVA bioconductor $2937$ $0.9969$ 0.001 080.001080.001\,080.001 08 $0.361$
+MAST bioconductor $2091$ $0.995$ 0.001 080.001080.001\,080.001 08 $0.361$
+FSA cran $499$ $0.9782$ 0.001 080.001080.001\,080.001 08 $0.361$
+Boruta cran $441$ $0.9754$ 0.001 080.001080.001\,080.001 08 $0.361$
+MixSIAR cran $377$ $0.9714$ 0.001 170.001170.001\,170.001 17 $0.4989$
+DRIMSeq bioconductor $335$ $0.9678$ 0.001 170.001170.001\,170.001 17 $0.4989$
+netDx bioconductor $174$ $0.9419$ 0.001 070.001070.001\,070.001 07 00
+COCOA bioconductor $162$ $0.9378$ 0.001 170.001170.001\,170.001 17 $0.4989$
+emoji pypi $161$ $0.9374$ 0.001 070.001070.001\,070.001 07 00
+aviso pypi $160$ $0.937$ 0.001 170.001170.001\,170.001 17 $0.4989$
+SiZer cran $159$ $0.9366$ 0.001 170.001170.001\,170.001 17 $0.4989$
+searchlight pypi $137$ $0.9273$ 0.001 070.001070.001\,070.001 07 00
+“Nebraska”
+vctrs cran $2$ $0.2516$ 0.024 380.024380.024\,380.024 38 $0.9928$
+withr cran $2$ $0.2516$ 0.024 380.024380.024\,380.024 38 $0.9928$
+isoband cran 00 00 0.024 380.024380.024\,380.024 38 $0.9928$
+gss cran $7$ $0.4991$ 0.012 120.012120.012\,120.012 12 $0.9849$
+SuppDists cran $3$ $0.3287$ 0.005 670.005670.005\,670.005 67 $0.9658$
+lfda cran $1$ $0.1232$ $0.0049$ $0.9581$
+ggtext cran $7$ $0.4991$ 0.004 820.004820.004\,820.004 82 $0.957$
+texttable pypi $1$ $0.1232$ 0.004 780.004780.004\,780.004 78 $0.9569$
+affyio bioconductor $5$ $0.4326$ 0.004 730.004730.004\,730.004 73 $0.956$
+dnaio pypi $1$ $0.1232$ 0.004 720.004720.004\,720.004 72 $0.9558$
+requests pypi $7$ $0.4991$ 0.004 190.004190.004\,190.004 19 $0.9499$
+tifffile pypi $4$ $0.3863$ 0.004 090.004090.004\,090.004 09 $0.9488$
+Table 4: Software quadrants, weighted graph (full)
+Software Ecosystem Mentions Centrality
+Count Percentile Value Percentile
+“Pasteur”
+PRISMA cran 15 7971579715\,79715 797 $0.9997$ 0.128 620.128620.128\,620.128 62 $1$
+DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.087 350.087350.087\,350.087 35 $0.9996$
+SAM cran 13 0481304813\,04813 048 $0.9992$ 0.105 450.105450.105\,450.105 45 $0.9999$
+limma bioconductor 13 4511345113\,45113 451 $0.9993$ 0.105 240.105240.105\,240.105 24 $0.9998$
+ggplot2 cran 14 4411444114\,44114 441 $0.9995$ 0.074 010.074010.074\,010.074 01 $0.9995$
+PERMANOVA cran 21 9602196021\,96021 960 $1$ 0.048 610.048610.048\,610.048 61 $0.999$
+edgeR bioconductor 12 9231292312\,92312 923 $0.9991$ 0.103 090.103090.103\,090.103 09 $0.9997$
+WGCNA cran 13 9151391513\,91513 915 $0.9994$ 0.051 950.051950.051\,950.051 95 $0.9992$
+pymol pypi 14 6041460414\,60414 604 $0.9996$ 0.046 610.046610.046\,610.046 61 $0.9989$
+tophat pypi $7642$ $0.9987$ 0.073 770.073770.073\,770.073 77 $0.9994$
+vegan cran 10 0121001210\,01210 012 $0.9989$ 0.051 720.051720.051\,720.051 72 $0.9991$
+lme4 cran 10 5101051010\,51010 510 $0.999$ 0.037 380.037380.037\,380.037 38 $0.9986$
+Popular
+GSVA bioconductor $2937$ $0.9969$ 0.001 360.001360.001\,360.001 36 $0.1214$
+MAST bioconductor $2091$ $0.995$ 0.001 360.001360.001\,360.001 36 $0.1214$
+GSA cran $844$ $0.9868$ 0.001 360.001360.001\,360.001 36 $0.2163$
+FSA cran $499$ $0.9782$ 0.001 360.001360.001\,360.001 36 $0.1508$
+Boruta cran $441$ $0.9754$ 0.001 360.001360.001\,360.001 36 $0.0968$
+MixSIAR cran $377$ $0.9714$ 0.001 380.001380.001\,380.001 38 $0.4265$
+AntWeb cran $282$ $0.9616$ 0.001 380.001380.001\,380.001 38 $0.4265$
+emoji pypi $161$ $0.9374$ 0.001 380.001380.001\,380.001 38 $0.4356$
+eureqa pypi $149$ $0.9331$ 0.001 380.001380.001\,380.001 38 $0.4265$
+searchlight pypi $137$ $0.9273$ 0.001 360.001360.001\,360.001 36 $0.259$
+pathfindR cran $133$ $0.9259$ 0.001 370.001370.001\,370.001 37 $0.4023$
+ChIPXpress bioconductor $125$ $0.9213$ 0.001 380.001380.001\,380.001 38 $0.4129$
+pypet pypi $125$ $0.9213$ 0.001 360.001360.001\,360.001 36 00
+ADAPTS cran $125$ $0.9213$ 0.001 370.001370.001\,370.001 37 $0.3452$
+“Nebraska”
+calculate_expression pypi $3$ $0.3287$ 0.018 720.018720.018\,720.018 72 $0.9956$
+zscore pypi $5$ $0.4326$ 0.008 510.008510.008\,510.008 51 $0.9893$
+webpages pypi $2$ $0.2516$ 0.007 070.007070.007\,070.007 07 $0.9872$
+setuptools pypi $6$ $0.4689$ 0.005 720.005720.005\,720.005 72 $0.9825$
+gpviz pypi $1$ $0.1232$ 0.004 530.004530.004\,530.004 53 $0.9759$
+glib pypi $3$ $0.3287$ 0.003 970.003970.003\,970.003 97 $0.9711$
+ego pypi $5$ $0.4326$ 0.003 260.003260.003\,260.003 26 $0.9607$
+xtick pypi $1$ $0.1232$ 0.003 090.003090.003\,090.003 09 $0.9575$
+pauvre pypi $5$ $0.4326$ 0.002 920.002920.002\,920.002 92 $0.9531$
+mappy pypi $4$ $0.3863$ 0.002 920.002920.002\,920.002 92 $0.9531$
+rle cran $2$ $0.2516$ 0.002 870.002870.002\,870.002 87 $0.9522$
+LineagePulse bioconductor $1$ $0.1232$ $0.0027$ $0.9474$
+Table 5: Software quadrants, weighted graph (LCC component only)
+Software Ecosystem Mentions Centrality
+Count Percentile Value Percentile
+“Pasteur”
+PRISMA cran 15 7971579715\,79715 797 $0.9996$ 0.257 560.257560.257\,560.257 56 $1$
+DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.174 920.174920.174\,920.174 92 $0.9995$
+SAM cran 13 0481304813\,04813 048 $0.999$ 0.211 160.211160.211\,160.211 16 $0.9999$
+limma bioconductor 13 4511345113\,45113 451 $0.9991$ 0.210 740.210740.210\,740.210 74 $0.9998$
+ggplot2 cran 14 4411444114\,44114 441 $0.9994$ $0.1482$ $0.9994$
+PERMANOVA cran 21 9602196021\,96021 960 $1$ 0.097 350.097350.097\,350.097 35 $0.9988$
+edgeR bioconductor 12 9231292312\,92312 923 $0.9989$ 0.206 430.206430.206\,430.206 43 $0.9996$
+WGCNA cran 13 9151391513\,91513 915 $0.9993$ 0.104 020.104020.104\,020.104 02 $0.999$
+pymol pypi 14 6041460414\,60414 604 $0.9995$ 0.093 330.093330.093\,330.093 33 $0.9986$
+tophat pypi $7642$ $0.9984$ 0.147 730.147730.147\,730.147 73 $0.9993$
+vegan cran 10 0121001210\,01210 012 $0.9986$ 0.103 570.103570.103\,570.103 57 $0.9989$
+lme4 cran 10 5101051010\,51010 510 $0.9988$ 0.074 850.074850.074\,850.074 85 $0.9983$
+Popular
+GSVA bioconductor $2937$ $0.996$ 0.002 720.002720.002\,720.002 72 $0.1185$
+MAST bioconductor $2091$ $0.9937$ 0.002 720.002720.002\,720.002 72 $0.1185$
+GSA cran $844$ $0.9833$ 0.002 720.002720.002\,720.002 72 $0.2167$
+FSA cran $499$ $0.9724$ 0.002 720.002720.002\,720.002 72 $0.148$
+Boruta cran $441$ $0.969$ 0.002 720.002720.002\,720.002 72 $0.0923$
+MixSIAR cran $377$ $0.9639$ 0.002 760.002760.002\,760.002 76 $0.3878$
+AntWeb cran $282$ $0.9516$ 0.002 760.002760.002\,760.002 76 $0.3878$
+emoji pypi $161$ $0.9215$ 0.002 760.002760.002\,760.002 76 $0.397$
+searchlight pypi $137$ $0.9091$ 0.002 730.002730.002\,730.002 73 $0.254$
+pathfindR cran $133$ $0.9073$ 0.002 750.002750.002\,750.002 75 $0.3677$
+ChIPXpress bioconductor $125$ $0.9018$ 0.002 760.002760.002\,760.002 76 $0.3765$
+pypet pypi $125$ $0.9018$ 0.002 720.002720.002\,720.002 72 00
+“Nebraska”
+calculate_expression pypi $3$ $0.2982$ $0.0375$ $0.9944$
+zscore pypi $5$ $0.3883$ 0.017 050.017050.017\,050.017 05 $0.9865$
+setuptools pypi $6$ $0.4215$ 0.011 460.011460.011\,460.011 46 $0.978$
+newick pypi $9$ $0.4957$ 0.008 060.008060.008\,060.008 06 $0.9649$
+xtick pypi $1$ $0.1305$ 0.006 180.006180.006\,180.006 18 $0.9469$
+nanolyse pypi $9$ $0.4957$ 0.005 950.005950.005\,950.005 95 $0.9443$
+splatter pypi $8$ $0.4742$ 0.005 840.005840.005\,840.005 84 $0.9422$
+pauvre pypi $5$ $0.3883$ 0.005 840.005840.005\,840.005 84 $0.9413$
+mappy pypi $4$ $0.3474$ 0.005 840.005840.005\,840.005 84 $0.9413$
+rle cran $2$ $0.232$ 0.005 740.005740.005\,740.005 74 $0.9402$
+scone pypi $8$ $0.4742$ 0.005 740.005740.005\,740.005 74 $0.9398$
+LineagePulse bioconductor $1$ $0.1305$ $0.0054$ $0.9344$
 4 Discussion
+Generally, the results verify our intuition that the majority of software packages will have a low Katz centrality and relatively few mentions.
+The outliers are then the most interesting packages, because they are central and often mentioned (“Pasteur”), or because their centrality outweighs their mentions (“Nebraska”), or vice versa (popular packages).
 The absence of cycles in the network of open-source software packages used in our sample of biomedical science is quite interesting.
 It suggests a more robust design structure in the universe of scientific software than in the general software world.
-The lack of cycles also improves our analysis with Katz centrality, as it excludes feedback mechanisms that can artificially inflate the centrality of packages on a loop.
+The lack of cycles also improves our analysis with Katz centrality, as the absence of loops excludes feedback mechanisms that can artificially inflate the centrality of packages on a loop.
 Packages with low mentions and high Katz centrality are therefore even more critical, as they represent essential dependencies of packages that enable large volumes of research without receiving direct attention.
-These are exactly the “Nebraskan” packages referenced by Munroe (2020) that we set out to find.
-Figure 3 seems to show one clear example of such a package: velvet, a PyPI package with a mention count of $197$, but a Katz centrality value calculated as $0.105$ for the largest connected cluster and as $0.053$ for the complete graph.
-Interestingly, for the unweighted representation, this package has a low centrality of $0.007$.
-Generally, the results verify our intuition that the majority of software packages will have a low Katz centrality and relatively few mentions.
-The outliers are then the most interesting packages, because they are central and often mentioned (“Pasteur”), or because their centrality outweighs their mentions (“Nebraska”), or vice versa (popular packages).
-The difference between the unweighted and weighted version of our analysis, in Figure 3 can help us parse different subtypes of these outliers.
-For example, the velvet package discussed above is central solely because it is mentioned by one highly cited paper (and dozens of less cited ones).
+These are the “Nebraskan” packages referenced by [1] that we set out to find.
+Figure 4 and Tables 3, 4, and 5 provide examples of these packages.
+As expected, most of these software packages are libraries used by other packages: vctrs [33] (a package for type manipulation in R), withr [34] (a set of tools for safely running functions that change global variables), isoband [35] (a package to generate isolines from gridded data), pauvre [36] (a package for plotting Oxford Nanopore data), newick [37] (read and write Newick data formatted files), setuptools [38] (a packaging library for Python projects).
+It is interesting that the prominent R “Nebraska” packages belong to the tidyverse system [39], and thus are co-authored by Hadley Wickham.
+The difference between the unweighted and weighted versions of our analysis, in Figure 4 can help us parse different subtypes of these outliers.
+For example, the velvet [40] package is central solely because it is mentioned by one highly cited paper (and dozens of less highly cited ones).
 This strategy is not central in the unweighted version of the analysis, since citation counts are ignored.
-Central packages under this representation therefore also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
-Altogether, our paper tries to quantify impact as importance, and to put some numbers behind the intuition.
+Central packages under this representation, therefore, also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
+It is interesting that cran and bioconductor dominate the “Nebraska” quadrant for the unweighted graph, while for the weighted graph, pypi is overrepresented.
+The reason for this difference may warrant additional research.
+Altogether, our paper tries to quantify impact as importance and to provide a set of metrics behind the intuition.
 Of course, this is by itself a perilous process: there could be various flavors of both impact and importance, and people may reasonably disagree about the details.
 Our contribution is to propose a certain quantification, which agrees with our intuitive understanding.
-Besides the possible problems with the exact definition of importance, our work has a number of limitations:
+In addition to possible problems with the exact definition of importance, our work has a number of limitations.
   1. 1.
-    The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions Dataset (see the discussion in (Istrate et al., 2022a)).
-Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms (Druskat et al., 2024).
-Notably, this affects the above-mentioned “Nebraskan” package velvet, which is linked to a Python package of the same name that provides signal processing and communications algorithms in Python (available on PyPI [^1] https://pypi.org/project/velvet/).
-Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short read assembly using de Bruijn graphs in genomics, as described in (Zerbino and Birney, 2008) and implemented mostly in C (Zerbino et al., 2014).
-Sometimes an open source package shares a name with a proprietary software (e.
-g.
-PRISM, PACE), which leads to confusion and incorrect usage counts.
+    The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions dataset (see the discussion in [11]).
+Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms [7].
+Notably, this affects the package velvet (see Figure 3a), which has been falsely linked to a Python package that provides signal processing and communications algorithms in Python [41].
+Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short-read assembly using de Bruijn graphs in genomics, as described in [42] and implemented mostly in C [40].
+While this does not change its centrality and our interpretation, the linking error led to its mislabelling as a PyPI package for the visualization in Figure 3.
+    Sometimes an open source package also shares a name with a proprietary software (e. g. PRISM, PACE), which leads to confusion and incorrect mention counts.
   2. 2.
-    Our approach is limited to the open source packages included into one of the chosen package managers.
-This excludes proprietary software like Excel or GraphPad Prism, and open source non-packaged software like Gephi (Bastian et al., 2009).
-The number of missing packages may be estimated from Table 14 of (Istrate et al., 2022a) which gives the coverage by GitHub at 64.39%, the coverage by CRAN at 8.36%, the coverage by PyPI at 5.86%, and the coverage by Bioconductor at 3.23%.
-GitHub coverage can be used as a proxy for the share of the open source software: while some packages – perhaps growing in number – are not hosted by GitHub, their fraction remains anecdotally low.
-If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e.
-between a quarter and one third of the open source packages.
+    Our approach is limited to the open-source packages available in one of the chosen ecosystems.
+This excludes proprietary software like Excel or GraphPad Prism, and open-source non-package software like Gephi [43] (a complied application).
+The number of missing packages may be estimated from Table 14 of [11], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%.
+GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [44].
+If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages.
 It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems.
-For open source non-package software, it may be feasible to identify source repository URLs and then to identify packages depended on by the software.
+For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software.
 That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools.
-On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services); nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT).
+On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services).
+Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT).
 In the long term, there is a possibility that the requirement to provide SBOMs for U.S.
-federal government purchasers222as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies.
+federal government purchasers111as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [45].
   3. 3.
-    In our version of the dependency chart, we always used the latest version of any package.
+    In our dependency network, we always used the latest version of any package.
 In reality, dependencies change between the versions.
@@
   4. 4.
-    Our method captures the off the shelf software packages used for the papers.
+    Our method captures the off-the-shelf software packages used for each article.
 It does not capture ad hoc software written specifically for the given paper, which may be contained in the code accompanying the papers.
-This code may load software libraries, and thus add software dependencies absent in our graph.
+This code may load software libraries and thus add software dependencies that are absent from our graph.
   5. 5.
-    The dependencies reflected in the package managers do not necessarily reflect the actual usage of the software.
-A library may be loaded, but not used for the actual execution path used in the given paper.
+    The dependencies reflected in the package metadata do not necessarily reflect the actual usage of the software.
+A library may be loaded, but not used for the actual execution by a given paper.
 A more reliable method would be to reproduce the analysis done in each paper and capture the actual library calls—which is probably prohibitively difficult for the number of papers covered.
@@
     Many Python and R packages ultimately depend on libraries outside their ecosystems (most often C and Fortran libraries).
-These dependencies are not easily captured by the package management software, and thus are outside of our analysis.
+These dependencies are not easily captured by the package management infrastructure and thus are outside of our analysis.
   7. 7.
-    Our dependency graph does not include the concept of alternatives, when package $A$ depends on the functionality that can be provided either by package $B_{1}$, or package $B_{2}$.
+    Our dependency network does not include the concept of alternatives, when package $A$ depends on the functionality that can be provided either by package $B_{1}$, or package $B_{2}$.
 Some package managers like Debian’s apt have the possibility of specifying alternatives, but this is not a common feature of package managers.
@@
     Different communities place different emphasis on infrastructure elements such as tests.
-This may lead to the over-representation of testing infrastructure packages in some cases, for example in the Python ecosystem.
-We do not try to attempt to decide which elements of the infrastructure are “actually important”, instead relying on the judgment of the community of users and developers—while recognizing that such judgment may be different across the landscape.
+This may lead to the over-representation of testing infrastructure packages in some cases, for example, in the Python ecosystem.
+We do not try to attempt to decide which elements of the infrastructure are “actually important”; instead, we rely on the judgment of the community of users and developers while recognizing that such judgment may be different across the landscape.
   9. 9.
-    There are several different Python package managers with greatly overlapping dependency graphs. We used PyPI for this study. Some alternatives like pip and Conda are also worth investigating.
+    There are different software package repositories and ecosystems for Python and R with greatly overlapping dependency graphs.
+We used PyPI, CRAN and Bioconductor for this study.
+Alternative repositories that provide packages for Python and R, such as Conda channels (conda-forge, bioconda, etc.), are also worth investigating.
 5 Conclusions and future work
-In this work, we investigated the network of software packages used in the biomedical papers available through PubMed Central.
+In this work, we investigated the network of software packages used in the biomedical literature available through PubMed Central.
 We demonstrated that these packages follow the structure of “Stokes’ diagram” (famous for ”Pasteur’s quadrant”), with some packages highly visible to the end users, and some packages less visible, but important in the network due to their dependencies.
 We discussed the use of Katz centrality in discovering the important packages and found examples of such packages in the biomedical field.
-We also discovered that the structure of the network of the software packages used in biomedical science is different from the structure of generic software packages: the absence of dependence loops may indicate a more robust, or intentional, design.
-These findings and insights as well as the underlying methodology can be used in funding and development of open source scientific software.
+These findings and insights, as well as the underlying methodology, can be used to identify critical, but low-visibility, open-source scientific software in need of targeted funding and support.
 Of course, there are many ways to extend this work.
-It would be interesting, for instance, to analyze common workflows for different disciplines, perhaps using co-occurrences of mentions, and map them into the dependency graph.
+It would be interesting, for instance, to analyze common workflows for different disciplines, perhaps using co-occurrences of mentions, and map them into the dependency network.
 This might help to discover packages important for specific sub-fields of biomedical sciences.
 Similarly, adding temporal dependencies to our graph may help to discover and predict development trends.
-Also, individual dependencies may differ in the actual impact that they have on the dependent software.
+Additionally, individual dependencies may differ in the actual impact that they have on the dependent software (i.e., a dependency that is only minimally utilized by the downstream software itself may be given less weight).
 Developing and applying a metric that describes these differences in impact would make it possible to refine our approach to quantifying the importance of research software dependencies.
@@
 acknowledges support from CNPq and FAPERJ (E-26/210.478/2024).
+J.H.
+acknowledges that this material is based upon work supported by the US National Science Foundation under Grant Nos.
+SMA- 1064209 (SciSIP), OCI-0943168 (VOSS) and ACI-1453548 (CAREER) as well as support from the Alfred P.
+Sloan Foundation, Grant/Award Number: 2016-7209 for the creation of the SoftCite dataset.
+The Chan Zuckerberg Initiative and the acknowledged funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
 Data and code availability
-The code for this paper is available at (Brown et al., 2023), and the data are available at (Brown, 2023).
+The code for this paper is available at [21], and the data are available at [24].
+Author contributions
+All authors contributed equally to the conceptualization, data curation, formal analysis, investigation, methodology, software, visualization, and writing (original draft, review & editing) of this work.

### v4

@@
 This lack of infrastructural understanding means we are limited in our ability to reward developers and maintainers, encourage collaboration and coordination, and direct science funding in a well-informed manner.
-Fig 1: XKCD on software dependency and ecosystem fragility [1].
-Also has roll-over text of Someday ImageMagick will finally break for good and we’ll have a long period of scrambling as we try to reassemble civilization from the rubble.
-Used with permission: https://creativecommons.org/licenses/by-nc/2.5/
 Scientific software is often invisible in publications, because citation practices in science have not changed at the same pace as software has become crucial.
@@
 These latter packages thus become indirect, transitive dependencies of the user-facing program.
-This complex nature of the network of dependencies has a number of implications, including for security [13] and computational reproducibility [14].
+This complex nature of the network of dependencies has a number of implications, including for security [Hatta2022Nebraska, 13] and computational reproducibility [14].
 In particular, all the work undertaken to develop, maintain, test, and distribute the underlying software is not directly visible in the publication record itself, and is not included in derivatives such as citation networks and knowledge graphs.
-The situation resembles the famous XKCD cartoon shown in  Figure 1, where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1].
+The situation resembles the famous XKCD cartoon where “all modern infrastructure” critically depends on “a project some random person in Nebraska has been thanklessly maintaining since 2003” [1].
 The word “thanklessly” is important in this context: being unknown, these critical pieces of software get much less recognition and credit than they deserve—and than the science needs.
 The absence of recognition may lead to dire consequences, for example, if, due to the lack of funding, the underlying libraries become vectors for malware attacks.
-There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termedtransitive credit [15, 16] but so far the problem has not yet been solved.
+There have been proposals to assign credit to these packages using the dependency structure to calculate what was been termed transitive credit [15, 16] but so far, the problem has not yet been solved.
 To implement the proposed measure, software projects must start to publicize the packages they rely on in a way that enables recognition and citation, beyond the technical dependency already recorded in manifest files such as pyproject.toml, DESCRIPTION, cargo.toml, pom.xml, etc.
@@
 Table 1: Summary statistics for the number of mentions in papers.
-“Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Gini” is Gini coefficient for the mentions count.
+“Number” is the number of packages, “Dependency only” is the fraction of packages not mentioned in any papers and pulled by the dependencies only, “Median” is the median number of mentions, “IQR” is the interquartile range for mentions count, “Max” is the maximal number of mentions, and “Gini” is the Gini coefficient for the mentions count.
 Ecosystem Number Dependency only Mentions count
@@
 pypi $5596$ 0.143 495 353 824 160.143495353824160.143\,495\,353\,824\,160.143 495 353 824 16 $4$ $17$ 20 3222032220\,32220 322 0.888 156 736 619 3960.8881567366193960.888\,156\,736\,619\,3960.888 156 736 619 396
-Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, “Gini” is Gini coefficient.
+Table 2: Summary statistics for Katz centrality. “Median” is the median centrality, “IQR” is the interquartile range, “Max” is the maximal centrality, and “Gini” is the Gini coefficient.
 Ecosystem Median IQR Max Gini
@@
 The distribution of centrality (Table 2) is smoother, especially for the weighted graph, but is still rather unequal.
-Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLoS CompBio for suggesting this metric).
+Perhaps an even more robust measure of “Nebraska” properties of the ecosystem is the number of packages that are not mentioned in any paper, but are in the dependencies of the packages that are mentioned (we are grateful to the anonymous reviewer of PLOS Computational Biology for suggesting this metric).
 It is shown in Table 1.
-In Tables 3, 4, and 5 we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
+In Tables 3, 4, and 5, we show the packages in “Pasteur”, Popular, and “Nebraska” packages using the following criteria:
   1. 1.
@@
 For example, the velvet [40] package is central solely because it is mentioned by one highly cited paper (and dozens of less highly cited ones).
-This strategy is not central in the unweighted version of the analysis, since citation counts are ignored.
+Such characteristics do not make a package central in the unweighted version of the analysis, since citation counts are ignored there.
 Central packages under this representation, therefore, also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
@@
 For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software.
-That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools.
+That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools [Hatta2022Nebraska].
 On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services).